You have 2 mixtures of a saline solution that are 12% by volume, and 20% by volume. You need 6 gallons of a solution that is 18% by volume for an experiment. How much of the 12% and the 20% solutions should you mix together?
Answer by mananth(16946)

Solution I --12%---x gal
Solution II--20%---(6-x)gal
Mixture total ---- 18% -----6 gal
12%*x+(6-x)*20%= 6*18%
multiply equation by 100
12x+(6-x)*20 = 108
12x+120-20x=108
12x-20x=108-120
-8x=-12x
/-8
x=-12/8
x=3/4 gallons --- 6% solution
Balance of 6 gallons is 20% solution
6-3/4
= 5 1/4 gallons